THE MAGELLANIC CLOUDS

NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 105 — 1 June 2010

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Figure 1: VLT-MAD superb resolution near-infrared image of the core of 30 Doradus (MNRAS, 405, 421 (2010)).

Editorial

Dear Colleagues,

It is my pleasure to present you the 105^{th} issue of the Magellanic Clouds Newsletter. It is particularly voluminous, in part (but certainly not only) due to the Herschel special issue of A&A and a number of OGLE papers. Well done all! Thanks especially to one person, who did not mind entering all 30 affiliations in the webform, in full.

Congratulations to Brian Williams for his beautiful Ph.D. thesis on supernova remnants! We wish him all the best in his further career.

The next issue is planned to be distributed on the 1^{st} of August 2010. Front cover pictures remain welcome!

Editorially Yours, Jacco van Loon

Refereed Journal Papers

A very young component in the pre-eminent starburst region of the Small Magellanic Cloud

M. Heydari-Malayeri^{1} and R. Selier^{1}

1LERMA, Observatoire de Paris, France

We present a study of the compact H ii region N 66A in the SMC pre-eminent starburst region N 66/NGC 346. Despite extensive research on various components of the N 66/NGC 346 complex, few studies have so far focused on N 66A, which is a special object in the whole complex and therefore deserves scrutiny. The study of this compact H ii region and its fellow objects seems important in the framework of massive star formation in the Magellanic Clouds. This analysis is mainly based on our ESO NTT observations, both imaging and spectroscopy, coupled with archive HST ACS data and Spitzer IRAC data. We derive a number of physical characteristics of the compact H ii region N 66A. Moreover, we present the spectral classification of the main exciting star of N 66A for the first time using spectroscopy. Its spectral features indicate a main sequence massive star of type O8. We compare this result with that based on the stellar Lyman continuum flux estimated from the ionized gas Hβ flux. The compact H ii region belongs to a rare class of H ii regions in the Magellanic Clouds, called High-Excitation Blobs (HEBs). N 66A most probably represents a very young massive star formation event in the N 66 complex, which has a range of ages.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1004.0657

The Optical Gravitational Lensing Experiment. The OGLE-III Catalog of Variable Stars. VII. Classical Cepheids in the Small Magellanic Cloud

I. Soszyński^{1}, R. Poleski^{1}, A. Udalski^{1}, M.K. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, Ł. Wyrzykowski^{3}, O. Szewczyk^{2} and K. Ulaczyk^{1}

1Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2Universidad de Concepción, Departamento de Astronomia, Casilla 160-C, Concepción, Chile 3Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK

The seventh part of the OGLE-III Catalog of Variable Stars (OIII-CVS) consists of 4630 classical Cepheids in the Small Magellanic Cloud (SMC). The sample includes 2626 fundamental-mode (F), 1644 first-overtone (1O), 83 secondovertone (2O), 59 double-mode F/1O, 215 double-mode 1O/2O, and three triple-mode classical Cepheids. For each object basic parameters, multi-epoch VI photometry collected within 8 or 13 years of observations, and finding charts are provided in the OGLE Internet archive. We present objects of particular interest: exceptionally numerous sample of single-mode second-overtone pulsators, five double Cepheids, two Cepheids with eclipsing variations superimposed on the pulsation light curves. At least 139 first-overtone Cepheids exhibit low-amplitude secondary variations with periods in the range 0.60–0.65 of the primary ones. These stars populate three distinct sequences in the Petersen diagram. The origin of this secondary modulation is still unknown. Contrary to the Large Magellanic Cloud (LMC) we found only a few candidates for anomalous Cepheids in the SMC. This fact may be a clue for the explanation of the origin of the anomalous Cepheids. The period and luminosity distributions of Cepheids in both Magellanic Clouds suggest that there are two or three populations of classical Cepheids in each of the galaxies. The main difference between the LMC and SMC lays in different numbers of Cepheids in each group. We fit the period–luminosity (PL) relations of SMC Cepheids and compare them with the LMC PL laws.

Published in Acta Astronomica (2010), 60, 17 Available from arXiv:1003.4518 and from http://acta.astrouw.edu.pl/Vol60/n1/pdf/pap 60 1 2.pdf

The SAGE-Spec Spitzer Legacy program: The life-cycle of dust and gas in the Large Magellanic Cloud

F. Kemper^{1}, Paul M. Woods^{1}, V. Antoniou^{2}, J.-P. Bernard^{3}, R.D. Blum^{4}, M.L. Boyer^{5}, J. Chan^{6}, C.-H.R. Chen^{7},

M. Cohen^{8}, C. Dijkstra^{9}, C. Engelbracht^{10}, M. Galametz^{11}, F. Galliano^{11}, C. Gielen^{12}, Karl D. Gordon^{5}, V. Gorjian^{13}, J. Harris^{10}, S. Hony^{11}, J.L. Hora^{14}, R. Indebetouw^{7,15}, O. Jones^{1}, A. Kawamura^{16}, E. Lagadec^{17,1}, B. Lawton^{5}, J.M. Leisenring^{7}, S.C. Madden^{11}, M. Marengo^{2,14}, M. Matsuura^{18,19}, I. McDonald^{1}, C. McGuire^{1}, M. Meixner^{5}, A.J. Mulia^{6}, B. O’Halloran^{20}, J.M. Oliveira^{21}, R. Paladini^{22}, D. Paradis^{22}, W.T. Reach^{22}, D. Rubin^{11},

K. Sandstrom^{8,23}, B.A. Sargent^{5}, M. Sewiło^{5}, B. Shiao^{5}, G.C. Sloan^{24}, A.K. Speck^{6}, S. Srinivasan^{25,26}, R. Szczerba^{27}, A.G.G.M. Tielens^{28}, E. van Aarle^{12}, S.D. Van Dyk^{22}, J.Th. van Loon^{21}, H. Van Winckel^{12}, Uma P.

Vijh^{29}, K. Volk^{5}, B.A. Whitney^{30}, A.N. Wilkins^{24} and A.A. Zijlstra^{1}

1Jodrell Bank Centre for Astrophysics, Alan Turing Building, School of Physics and Astronomy, The University of Manchester, Oxford Road, Manchester, M13 9PL, UK 2Department of Physics and Astronomy, Iowa State University, Ames, IA 50011, USA 3Centre d’Étude Spatiale des Rayonnements, 9 Av. du Colonel Roche, BP 44346, 31028 Toulouse cedex 4, France 4NOAO, 950 North Cherry Avenue, Tucson, AZ 85719, USA 5Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 6Physics & Astronomy Department, University of Missouri, Columbia, MO 65211, USA 7Department of Astronomy, University of Virginia, P.O. Box 400325, Charlottesville, VA 22904, USA 8Radio Astronomy Laboratory, University of California at Berkeley, 601 Campbell Hall, Berkeley, CA 94720-3411, USA 9Passiebloemweg 31, 1338 TT Almere, The Netherlands 10Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721, USA 11Laboratoire AIM, CEA/DSM–CNRS–Université Paris Diderot DAPNIA/Service d’Astrophysique Bât. 709, CEA-Saclay F-91191 Gifsur-Yvette Cédex, France 12Instituut Voor Sterrenkunde, KU Leuven, Celestijnenlaan 200D, 3001 Leuven (Heverlee), Belgium 13JPL/Caltech, MS 169-506, 4800 Oak grove Dr., Pasadena, CA 91109, USA 14Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, MS 65, Cambridge, MA 02138-1516, USA 15National Radio Astronomy Observatory, 520 Edgemont Road, Charlottesville, VA 22903, USA 16Department of Astrophysics, Nagoya University, Chikusa-Ku, Nagoya 464-01, Japan 17ESO Headquarters Garching, Karl-Schwarzschild-Str. 2, D-85748 Garching bei München, Germany 18Institute of Origins, Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 19Institute of Origins, Mullard Space Science Laboratory, University College London, Holmbury St. Mary, Dorking, Surrey RH5 6NT, UK 20Astrophysics Group, Imperial College London, Blackett Laboratory, Prince Consort Road, London, SW7 2AZ,UK 21School of Physical & Geographical Sciences, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK 22Spitzer Science Center, California Institute of Technology, MS 220-6, Pasadena, CA 91125, USA 23Max-Planck-Institut für Astronomie, D-69117 Heidelberg, Germany 24Department of Astronomy, Cornell University, Ithaca, NY 14853-6801, USA 25Department of Physics and Astronomy, Johns Hopkins University, Homewood Campus, Baltimore, MD 21218, USA 26Institut d’Astrophysique de Paris, 98 bis, Boulevard Arago, Paris 75014, France 27N. Copernicus Astronomical Center, Rabianska 8, 87-100 Torun, Poland 28Leiden Observatory, P.O. Box 9513, NL-2300 RA Leiden, The Netherlands 29Ritter Astrophysical Research Center, University of Toledo, Toledo OH 43606, USA 30Space Science Institute, 4750 Walnut Street, Suite 205, Boulder, CO 80301, USA

The SAGE-Spec Spitzer Legacy program is a spectroscopic follow-up to the SAGE-LMC photometric survey of the Large Magellanic Cloud carried out with the Spitzer Space Telescope. We present an overview of SAGE-Spec and some of its first results. The SAGE-Spec program aims to study the life cycle of gas and dust in the Large Magellanic Cloud, and to provide information essential to the classification of the point sources observed in the earlier SAGE-LMC photometric survey. We acquired 224.6 hours of observations using the InfraRed Spectrograph and the SED mode of the Multiband Imaging Photometer for Spitzer. The SAGE-Spec data, along with archival Spitzer spectroscopy of objects in the Large Magellanic Cloud, are reduced and delivered to the community. We discuss the observing strategy, the specific data reduction pipelines applied and the dissemination of data products to the scientific community. Initial science results include the first detection of an extragalactic “21 μm” feature towards an evolved star and elucidation of the nature of disks around RV Tauri stars in the Large Magellanic Cloud. Towards some young stars, ice features are observed in absorption. We also serendipitously observed a background quasar, at a redshift of z ≈ 0.14, which appears to be host-less.

Accepted for publication in PASP Available from arXiv:1004.1142

Concerning the distance to the center of the Milky Way and its structure

Daniel J. Majaess^{1,2}

1Saint Mary’s University, Halifax, Nova Scotia, Canada 2The Abbey Ridge Observatory, Stillwater Lake, Nova Scotia, Canada

The distance to the Galactic center inferred from OGLE RR Lyrae variables observed in the direction of the bulge is R_{0} = 8.1 ± 0.6 kpc. An accurate determination of R_{0} is hindered by countless effects that include an ambiguous extinction law, a bias for smaller values of R_{0} because of a preferential sampling of variable stars toward the near side of the bulge owing to extinction, and an uncertainty in characterizing how a mean distance to the group of variable stars relates to R_{0}. A VI-based period–reddening relation for RR Lyrae variables is derived to map extinction throughout the bulge. The reddening inferred from RR Lyrae variables in the Galactic Bulge, LMC, SMC, and IC 1613 match that established from OGLE red clump giants and classical Cepheids. RR Lyrae variables obey a period–colour (VI) relation that is relatively insensitive to metallicity. Edge-on and face-on illustrations of the Milky Way are constructed by mapping the bulge RR Lyrae variables in tandem with cataloged red clump giants, globular clusters, planetary nebulae, classical Cepheids, young open clusters, H ii regions, and molecular clouds. The sample of RR Lyrae variables do not trace a prominent Galactic bar or triaxial bulge oriented at ϕ ∼ 25^{∘}.

Accepted for publication in Acta A. Available from arXiv:1002.2743 and from http://acta.astrouw.edu.pl/Vol60/n1/a 60 1 4.html

Ages and luminosities of young SMC/LMC star clusters and the recent star formation history of the Clouds

Katharina Glatt ^{1,2}, Eva K. Grebel ^{1,2} and Andreas Koch^{3}

1Department of Physics and Astronomy, Klingelbergstr. 82, CH-4056 Basel, Switzerland 2Astronomisches Rechen-Institut, Zentrum für Astronomie der Universität Heidelberg, Mönchhofstr. 12–14, D-69120, Germany 3Department of Physics and Astronomy, University of Leicester, University Road, Leicester, LE1 7RH, UK

In this paper we discuss the age and spatial distribution of young (age < 1 Gyr) SMC and LMC clusters using data from the Magellanic Cloud Photometric Surveys. Luminosities are calculated for all age-dated clusters. Ages of 324 and 1193 populous star clusters in the Small and the Large Magellanic Cloud have been determined fitting Padova and Geneva isochrone models to their resolved color–magnitude diagrams. The clusters cover an age range between 10 Myr and 1 Gyr in each galaxy. For the SMC a constant distance modulus of (m − M)_{0} = 18.90 and a metallicity of Z = 0.004 were adopted. For the LMC, we used a constant distance modulus of (m − M)_{0} = 18.50 and a metallicity of Z = 0.008. For both galaxies, we used a variable color excess to derive the cluster ages. We find two periods of enhanced cluster formation in both galaxies at 160 Myr and 630 Myr (SMC) and at 125 Myr and 800 Myr (LMC). We present the spatially resolved recent star formation history of both Clouds based on young star clusters. The first peak may have been triggered by a close encounter between the SMC and the LMC. In both galaxies the youngest clusters reside in the supergiant shells, giant shells, the inter-shell regions, and toward regions with a high Hα content, suggesting that their formation is related to expansion and shell–shell interaction. Most of the clusters are older than the dynamical age of the supergiant shells. No evidence for cluster dissolution was found. Computed V-band luminosities show a trend for fainter magnitudes with increasing age as well as a trend for brighter magnitudes with increasing apparent cluster radii.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1004.1247

The Optical Gravitational Lensing Experiment. The OGLE-III Catalog of Variable Stars. VI. Delta Scuti stars in the Large Magellanic Cloud

R. Poleski^{1}, I. Soszyński^{1}, A. Udalski^{1}, M.K. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, Ł. Wyrzykowski^{3}, O. Szewczyk^{2} and K. Ulaczyk^{1}

1Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2Universidad de Concepción, Departamento de Fisica, Casilla 160-C, Concepción, Chile 3Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK

The sixth part of the OGLE-III Catalog of Variable Stars presents δ Sct pulsators in the Large Magellanic Cloud. Altogether 2786 variables were found and amongst them 92 are multi-mode objects, including 67 stars pulsating in the fundamental mode and the first overtone (F/1O), 9 double-mode pulsators with various combinations of the first three overtones excited (1O/2O, 2O/3O and 1O/3O pulsators), and two triple mode (F/1O/2O) δ Sct stars. In total 1490 of stars are marked as uncertain, due to scattered photometry and small amplitudes. For single-mode objects it was not possible to unambiguously identify pulsation mode, however we suggest the most of the single-mode variables pulsate in the first overtone.

Published in Acta Astronomica 60, 1 Available from arXiv:1004.0950 and from http://ogle.astrouw.edu.pl/cont/4 main/var/ogleiii/dsctlmc/text.html

Physical properties of giant molecular clouds in the Large Magellanic Cloud

A. Hughes^{1,2}, T. Wong^{3}, J. Ott^{4}, E. Muller^{5}, J.L. Pineda^{6}, Y. Mizuno^{5}, J.-P. Bernard^{7}, D. Paradis^{8}, S. Maddison^{1}, W.T. Reach^{8,9}, L. Staveley-Smith^{10}, A. Kawamura^{5}, M. Meixner^{11,12}, S. Kim^{13}, T. Onishi^{5,14}, N.

Mizuno^{15} and Y. Fukui^{5}

1Centre for Supercomputing and Astrophysics, Swinburne University of Technology, Hawthorn VIC 3122, Australia 2CSIRO Australia Telescope National Facility, P.O. Box 76, Epping NSW 1710, Australia 3Astronomy Department, University of Illinois, 1002 W. Green St., Urbana, IL 61801, USA 4National Radio Astronomy Observatory, P.O. Box O, 1003 Lopezville Road, Socorro, NM 87801, USA 5Department of Astrophysics, Nagoya University, Furo-cho, Chikusa-ku, Nagoya 464-8602, Japan 6Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA 91109-8099, USA 7Centre d’Étude Spatiale des Rayonnements, Universite Paul Sabatier, 9 Av. du Colonel Roche, BP 44346, Toulouse, France 8Spitzer Science Center, Caltech, MS220-6, Pasadena, CA 91125, USA 9Stratospheric Observatory for Infrared Astronomy, Universities Space Research Association, Mail Stop 211-3, Moffett Field, CA 94035, USA 10International Centre for Radio Astronomy Research, M468, University of Western Australia, 35 Stirling Hwy, Crawley, WA 6009, Australia 11Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 12Radio and Geoastronomy Division, Harvard-Smithsonian for Astrophysics, 60 Garden St., MS 42, Cambridge, MA 02138-1516, USA 13Department of Astronomy and Space Science, Sejong University, KwangJin-gu, KunJa-dong 98, Seoul 143-747, South Korea 14Department of Physical Science, Osaka Prefecture University, Gakeun 1-1, Sakai, Osaka 599-8531, Japan 15ALMA-J Project Office, National Astronomical Observatory of Japan, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan

The Magellanic Mopra Assessment (MAGMA) is a high angular resolution CO mapping survey of giant molecular clouds (GMCs) in the Large and Small Magellanic Clouds using the Mopra Telescope. Here we report on the basic physical properties of 125 GMCs in the LMC that have been surveyed to date. The observed clouds exhibit scaling relations that are similar to those determined for Galactic GMCs, although LMC clouds have narrower linewidths and lower CO luminosities than Galactic clouds of a similar size. The average mass surface density of the LMC clouds is 50 M_{⊙} pc^{−2}, approximately half that of GMCs in the inner Milky Way. We compare the properties of GMCs with and without signs of massive star formation, finding that non-star-forming GMCs have lower peak CO brightness than star-forming GMCs. We compare the properties of GMCs with estimates for local interstellar conditions: specifically, we investigate the H i column density, radiation field, stellar mass surface density and the external pressure. Very few cloud properties demonstrate a clear dependence on the environment; the exceptions are significant positive correlations between i) the H i column density and the GMC velocity dispersion, ii) the stellar mass surface density and the average peak CO brightness, and iii) the stellar mass surface density and the CO surface brightness. The molecular mass surface density of GMCs without signs of massive star formation shows no dependence on the local radiation field, which is inconsistent with the photoionization-regulated star formation theory proposed by McKee (1989). We find some evidence that the mass surface density of the MAGMA clouds increases with the interstellar pressure, as proposed by Elmegreen (1989), but the detailed predictions of this model are not fulfilled once estimates for the local radiation field, metallicity and GMC envelope mass are taken into account.

Accepted for publication in MNRAS Available from arXiv:1004.2094

Can massive Be/Oe stars be progenitors of long gamma ray bursts?

C. Martayan^{1,2}, J. Zorec^{3}, Y. Frémat^{4} and S. Ekström^{5}

1European Organization for Astronomical Research in the Southern Hemisphere, Alonso de Cordova 3107, Vitacura, Santiago de Chile, Chile 2GEPI, Observatoire de Paris, CNRS, Université Paris Diderot, 5 place Jules Janssen, 92195 Meudon Cedex, France 3Institut d’Astrophysique de Paris, UMR7095, CNRS, Université Marie & Pierre Curie, 98bis Boulevard Arago 75014 Paris, France 4Royal Observatory of Belgium, 3 avenue circulaire, 1180 Brussels, Belgium 5Geneva Observatory, University of Geneva, Maillettes 51, 1290 Sauverny, Switzerland

Context: The identification of long-gamma-ray-bursts (LGRBs) is still uncertain, although the collapsar engine of fast-rotating massive stars is gaining a strong consensus. Aims: We propose that low-metallicity Be and Oe stars, which are massive fast rotators, as potential LGRBs progenitors. Methods: We checked this hypothesis by 1) testing the global specific angular momentum of Oe/Be stars in the ZAMS with the SMC metallicity, 2) comparing the ZAMS (Ω/Ω_{c}, M/M_{⊙}) parameters of these stars with the area predicted theoretically for progenitors with metallicity Z = 0.002, and 3) calculating the expected rate of LGRBs/year/galaxy and comparing them with the observed ones. To this end, we determined the ZAMS linear and angular rotational velocities for SMC Be and Oe stars using the observed v sin i parameters, corrected from the underestimation induced by the gravitational darkening effect. Results: The angular velocities of SMC Oe/Be stars are on average ⟨Ω/Ω_{c}⟩ = 0.95 in the ZAMS. These velocities are in the area theoretically predicted for the LGRBs progenitors. We estimated the yearly rate per galaxy of LGRBs and the number of LGRBs produced in the local Universe up to z = 0.2. We have considered that the mass range of LGRB progenitors corresponds to stars hotter than spectral types B0–B1 and used individual beaming angles from 5 to 15∘. We thus obtain R_{LGRB}^{pred} ∼ 10^{−7} to ∼ 10^{−6} LGRBs/year/galaxy, which represents on average 2 to 14 LGRB predicted events in the local Universe during the past 11 years. The predicted rates could widely surpass the observed ones [(0.2–3)×10^{−7} LGRBs/year/galaxy; 8 LGRBs observed in the local Universe during the last 11 years] if the stellar counts were made from the spectral type B1–B2, in accordance with the expected apparent spectral types of the appropriate massive fast rotators. Conclusion: We conclude that the massive Be/Oe stars with SMC metallicity could be LGRBs progenitors. Nevertheless,other SMC O/B stars without emission lines, which have high enough specific angular momentum, can enhance the predicted R_{LGRB} rate.

Accepted for publication in A&A Available from arXiv:1004.3362

The Spitzer Survey of the Small Magellanic Cloud (S^{3}MC): Insights into the life-cycle of Polycyclic Aromatic Hydrocarbons

Karin M. Sandstrom^{1,2}, Alberto D. Bolatto^{3}, B.T. Draine^{4}, Caroline Bot^{5} and Snežana Stanimirović^{6}

1Astronomy Department, University of California, Berkeley, USA 2Max Planck Institut für Astronomie, Heidelberg, Germany 3Department of Astronomy and Laboratory for Millimeter-wave Astronomy, University of Maryland, College Park, USA 4Department of Astrophysical Sciences, Princeton University, USA 5Observatoire Astronomiques de Strasbourg, Université Louis Pasteur, France 6Astronomy Department, University of Wisconsin, Madison, USA

We present the results of modeling dust spectral energy distributions (SEDs) across the Small Magellanic Cloud (SMC) with the aim of mapping the distribution of polycyclic aromatic hydrocarbons (PAHs) in a low-metallicity environment. Using Spitzer Survey of the SMC (S^{3}MC) photometry from 3.6 to 160 μm over the main star-forming regions of the Wing and Bar of the SMC along with spectral mapping observations from 5 to 38 μm from the Spitzer Spectroscopic Survey of the Small Magellanic Cloud (S^{4}MC) in selected regions, we model the dust spectral energy distribution and emission spectrum to determine the fraction of dust in PAHs across the SMC. We use the regions of overlaping photometry and spectroscopy to test the reliability of the PAH fraction as determined from SED fits alone. The PAH fraction in the SMC is low compared to the Milky Way and variable — with relatively high fractions (q_{PAH} ∼ 1–2%) in molecular clouds and low fractions in the diffuse ISM (average ⟨q_{PAH}⟩ = 0.6%). We use the map of PAH fraction across the SMC to test a number of ideas regarding the production, destruction and processing of PAHs in the ISM. We find weak or no correlation between the PAH fraction and the distribution of carbon AGB stars, the location of supergiant H i shells and young supernova remnants, and the turbulent Mach number. We find that the PAH fraction is correlated with CO intensity, peaks in the dust surface density and the molecular gas surface density as determined from 160 μm emission. The PAH fraction is high in regions of active star-formation, as predicted by its correlation with molecular gas, but is supressed in H ii regions. Because the PAH fraction in the diffuse ISM is generally very low — in accordance with previous work on modeling the integrated SED of the SMC — and the PAH fraction is relatively high in molecular regions, we suggest that PAHs are destroyed in the diffuse ISM of the SMC and/or PAHs are forming in molecular clouds. We discuss the implications of these observations for our understanding of the PAH life cycle, particularly in low-metallicity and/or primordial galaxies.

Accepted for publication in The Astrophysical Journal Available from arXiv:1003.4516

Searching for diffuse nonthermal X-rays from the superbubbles N 11 and N 51D in the Large Magellanic Cloud

Hiroya Yamaguchi^{1}, Makoto Sawada^{2} and Aya Bamba^{3,4}

1RIKEN, Japan 2Kyoto University, Japan 3Dublin Institute for Advanced Studies, Republic of Ireland 4Institute of Space and Astronautical Science, Japan

We report on observations of the superbubbles (SBs) N 11 and N 51D in the Large Magellanic Cloud (LMC) with Suzaku and XMM-Newton. The interior of both SBs exhibits diffuse X-ray emission, which is well represented by thin thermal plasma models with a temperature of 0.2–0.3 keV. The presence of nonthermal emission, claimed in previous works, is much less evident in our careful investigation. The 3-σ upper limits of 2–10 keV flux are 3.6 × 10^{−14} ergs cm^{−2} s^{−1} and 4.7 × 10^{−14} ergs cm^{−2} s^{−1} for N 11 and N 51D, respectively. The previous claims of the detection of nonthermal emission are probably due to the inaccurate estimation of the non X-ray background. We conclude that no credible nonthermal emission has been detected from the SBs in the LMC, with the exception of 30 Dor C.

Published in Astrophysical Journal Available from arXiv:1004.0753

The OGLE view of microlensing towards the Magellanic Clouds. II. OGLE-II SMC data

Ł. Wyrzykowski^{1,2}, S. Kozłowski^{3}, J. Skowron^{2}, V. Belokurov^{1}, M.C. Smith^{1}, A. Udalski^{2}, M.K. Szymański^{2}, M. Kubiak^{2}, G. Pietrzyński^{2,4}, I. Soszyński^{2} and O. Szewczyk^{2,4}

1Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 2Warsaw University Astronomical Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 3Department of Astronomy, Ohio State University, 140 W. 18^{th} Ave., Columbus, OH 43210, USA 4Universidad de Concepción, Departamento de Fisica, Astronomy Group, Casilla 160-C, Concepción, Chile

The primary goal of this paper is to provide the evidence that can either prove or falsify the hypothesis that dark matter in the Galactic halo can clump into stellar-mass compact objects. If such objects existed, they would act as lenses to external sources in the Magellanic Clouds, giving rise to an observable effect of microlensing. We present the results of our search for such events, based on the data from the second phase of the OGLE survey (1996–2000) towards the SMC. The data set we used is comprised of 2.1 million monitored sources distributed over an area of 2.4 square degrees. We found only one microlensing event candidate, however its poor quality light curve limited our discussion on the exact distance to the lensing object. Given a single event, taking the blending (crowding of stars) into account for the detection efficiency simulations, and deriving the HST-corrected number of monitored stars, the microlensing optical depth is τ = (1.55±1.55)×10^{−7}. This result is consistent with the expected SMC self-lensing signal, with no need of introducing dark matter microlenses. Rejecting the unconvincing event leads to the upper limit on the fraction of dark matter in the form of MACHOs to f < 20 per cent for deflectors’ masses around 0.4 M_{⊙} and f < 11 per cent for masses between 0.003 and 0.2 M_{⊙} (95 per cent confidence limit). Our result indicates that the Milky Way’s dark matter is unlikely to be clumpy and form compact objects in the sub-solar-mass range.

Accepted for publication in MNRAS Available from arXiv:1004.5247 and from http://ogle.astrouw.edu.pl/

On the stellar content of Carina dwarf galaxy

Bono et al.^{1}

1Università di Roma ”Tor Vergata”, Italy

We present U,B,V,I photometry of the Carina dwarf spheroidal galaxy, based on more than 4,000 CCD images. Special attention was given to the photometric calibration, and the precision for the B,V,I bands is better than 0.01 mag. We compared in the V,B–V and V,B–I color–magnitude diagrams (CMDs) Carina with three Globular Clusters (GCs, M 53, M 55, M 79). We find that only the more metal-poor GCs (M 55, [Fe/H]= −1.85; M 53, [Fe/H]= −2.02 dex) provide a good match with the Carina giant branch. We performed a similar comparison in the V,V–I CMD with three SMC intermediate-age clusters (IACs, Kron 3, NGC 339, Lindsay 38). We find that the color extent of the SGB of the two more metal-rich IACs (Kron 3, [Fe/H]= −1.08; NGC 339, [Fe/H]= −1.36 dex) is smaller than the range among Carina’s intermediate-age stars. However, the ridge line of the more metal-poor IAC (Lindsay 38, [Fe/H]= −1.59 dex) agrees quite well with the Carina intermediate-age stars. These findings indicate that Carina’s old stellar population is metal-poor and seems to have a limited spread in metallicity (Δ[Fe/H]= 0.2–0.3 dex). Carina intermediate-age stars can hardly be more metal-rich than Lindsay 38 and its spread in metallicity appears modest. We also find that the synthetic CMD constructed assuming a metallicity spread of 0.5 dex for intermediate-age stars predicts evolutionary features not supported by observations. The above results are at odds with recent spectroscopic investigations suggesting that Carina stars cover a broad range in metallicity (Δ[Fe/H]∼ 1–2 dex). We present a new method to estimate the metallicity of complex stellar systems using the difference in color between the red clump and the middle of the RR Lyrae instability strip. The observed colors of Carina’s evolved stars indicate a metallicity of [Fe/H]= −1.70 ± 0.19 dex, which agrees quite well with spectroscopic measurements.

Accepted for publication in PASP Available from arXiv:1004.2559

Spitzer analysis of H ii region complexes in the Magellanic Clouds: Determining a suitable monochromatic obscured star formation indicator

B. Lawton^{1}, K.D. Gordon^{1}, B. Babler^{2}, M. Block^{3}, A.D. Bolatto^{4}, S. Bracker^{2}, L.R. Carlson^{5}, C.W. Engelbracht^{3}, J.L. Hora^{6}, R. Indebetouw^{7}, S.C. Madden^{8}, M. Meade^{2}, M. Meixner^{1}, K. Misselt^{3}, M.S. Oey^{9}, J.M. Oliveira^{10}, T.

Robitaille^{6}, M. Sewiło^{1}, B. Shiao^{1}, U.P. Vijh^{11} and B. Whitney^{12}

1Space Telescope Science Institute, Baltimore, MD, 21218, USA 2Department of Astronomy, University of Wisconsin-Madison, 475 N. Charter St., Madison, WI 53706, USA 3Steward Observatory, University of Arizona, 933 North Cherry Ave., Tucson, AZ 85721, USA 4Department of Astronomy, University of Maryland, College Park, MD 20742, USA 5Department of Physics and Astronomy, Johns Hopkins University, Baltimore, MD, USA 6Harvard-Smithsonian, CfA, 60 Garden St., MS 65, Cambridge, MA 02138-1516, USA 7Department of Astronomy, University of Virginia, P.O. Box 3818, Charlottesville, VA 22903-0818, USA 8Service d’Astrophysique, CEA/Saclay, l’Orme des Merisiers, 91191 Gifsur-Yvette, France 9Department of Astronomy, University of Michigan, Ann Arbor, MI 48109-1042, USA 10School of Physical and Geographical Sciences, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK 11Ritter Astrophysical Research Center, University of Toledo, Toledo, OH 43606, USA 12Space Science Institute, 4750 Walnut St., Suite 205, Boulder, CO 80301, USA

H ii regions are the birth places of stars, and as such they provide the best measure of current star formation rates (SFRs) in galaxies. The close proximity of the Magellanic Clouds allows us to probe the nature of these star forming regions at small spatial scales. To study the H ii regions, we compute the bolometric infrared flux, or total infrared (TIR), by integrating the flux from 8 to 500 μm. The TIR provides a measure of the obscured star formation because the UV photons from hot young stars are absorbed by dust and re-emitted across the mid-to-far-infrared (IR) spectrum. We aim to determine the monochromatic IR band that most accurately traces the TIR and produces an accurate obscured SFR over large spatial scales. We present the spatial analysis, via aperture/annulus photometry, of 16 Large Magellanic Cloud (LMC) and 16 Small Magellanic Cloud (SMC) H ii region complexes using the Spitzer Space Telescope’s IRAC (3.6, 4.5, 8 μm) and MIPS (24, 70, 160 μm) bands. Ultraviolet rocket data (1500 and 1900 Å) and SHASSA Hα data are also included. All data are convolved to the MIPS 160 μm resolution (40′′ full width at half-maximum), and apertures have a minimum radius of 35′′. The IRAC, MIPS, UV, and Hα spatial analysis are compared with the spatial analysis of the TIR. We find that nearly all of the LMC and SMC H ii region spectral energy distributions (SEDs) peak around 70 μm at all radii, from ∼ 10 to ∼ 400 pc from the central ionizing sources. As a result, we find the following: the sizes of H ii regions as probed by 70 μm is approximately equal to the sizes as probed by TIR (≈ 70 pc in radius); the radial profile of the 70 μm flux, normalized by TIR, is constant at all radii (70 μm∼ 0.45TIR); the 1σ standard deviation of the 70 μm fluxes, normalized by TIR, is a lower fraction of the mean (0.05–0.12 out to ∼ 220 pc) than the normalized 8, 24, and 160 μm normalized fluxes (0.12–0.52); and these results are the same for the LMC and the SMC. From these results, we argue that 70 μm is the most suitable IR band to use as a monochromatic obscured star formation indicator because it most accurately reproduces the TIR of H ii regions in the LMC and SMC and over large spatial scales. We also explore the general trends of the 8, 24, 70, and 160 μm bands in the LMC and SMC H ii region SEDs, radial surface brightness profiles, sizes, and normalized (by TIR) radial flux profiles. We derive an obscured SFR equation that is modified from the literature to use 70 μm luminosity, SFR(M_{⊙} yr^{−1}) = 9.7(0.7) × 10^{−44} L_{70} (ergs s^{−1}), which is applicable from 10 to 300 pc distance from the center of an H ii region. We include an analysis of the spatial variations around H ii regions between the obscured star formation indicators given by the IR and the unobscured star formation indicators given by UV and Hα. We compute obscured and unobscured SFRs using equations from the literature and examine the spatial variations of the SFRs around H ii regions.

Accepted for publication in Astrophysical Journal Available from arXiv:1005.0036

A massive runaway star from 30 Doradus

C.J. Evans^{1}, N.R. Walborn, P.A. Crowther, V. Hénault-Brunet, D. Massa, W.D. Taylor, I.D. Howarth, H. Sana, D.J. Lennon and J.Th. van Loon

1UK ATC, Scotland

We present the first ultraviolet (UV) and multi-epoch optical spectroscopy of 30 Dor 016, a massive O2-type star on the periphery of 30 Doradus in the Large Magellanic Cloud. The UV data were obtained with the Cosmic Origins Spectrograph on the Hubble Space Telescope as part of the Servicing Mission Observatory Verification program after Servicing Mission 4, and reveal #016 to have one of the fastest stellar winds known. From analysis of the C iv λλ1548– 51 doublet we find a terminal velocity, v_{∞} = 3450 ± 50 km s^{−1}. Optical spectroscopy is from the VLT–FLAMES Tarantula Survey, from which we rule out a massive companion (with 2 days < P < 1 yr) to a confidence of 98%. The radial velocity of #016 is offset from the systemic value by −85 km s^{−1}, suggesting that the star has traveled the 120 pc from the core of 30 Doradus as a runaway, ejected via dynamical interactions.

Published in ApJ, 715, L74 Available from arXiv:1004.5402

INTEGRAL deep observations of the Small Magellanic Cloud

M.J. Coe et al.^{1}

1University of Southampton, UK

Deep observations of the Small Magellanic Cloud (SMC) and region were carried out in the hard X-ray band by the INTEGRAL observatory in 2008–2009. The field of view of the instrument permitted simultaneous coverage of the entire SMC and the eastern end of the Magellanic Bridge. In total, INTEGRAL detected seven sources in the SMC and five in the Magellanic Bridge; the majority of the sources were previously unknown systems. Several of the new sources were detected undergoing bright X-ray outbursts and all the sources exhibited transient behaviour except the supergiant system SMC X-1. They are all thought to be High Mass X-ray Binary (HMXB) systems in which the compact object is a neutron star.

Accepted for publication in MNRAS Available from arXiv:1004.2219

Evolution of the Chandra CCD spectra of SNR 1987A: Probing the reflected-shock picture

Svetozar A. Zhekov^{1,4}, Sangwook Park^{2}, Richard McCray^{1}, Judith L. Racusin^{3} and David N. Burrows^{2}

1JILA, University of Colorado, Boulder, CO 80309-0440, USA 2Department of Astronomy and Astrophysics, Pennsylvania State University, 525 Davey Laboratory, University Park, PA 16802, USA 3NASA GSFC, 8800 Greenbelt Rd., Code 661, Greenbelt, MD 20771, USA 4On leave from Space Research Institute, Sofia, Bulgaria

We continue to explore the validity of the reflected shock structure (RSS) picture in SNR 1987A that was proposed in our previous analyses of the X-ray emission from this object. We used an improved version of our RSS model in a global analysis of 14 CCD spectra from the monitoring program with Chandra. In the framework of the RSS picture, we are able to match both the expansion velocity curve deduced from the analysis of the X-ray images and light curve. Using a simplified analysis, we also show that the X-rays and the non-thermal radio emission may originate from the same shock structure (the blast wave). We believe that using the RSS model in the analysis of grating data from the Chandra monitoring program of SNR 1987A that cover a long enough time interval, will allow us to build a more realistic physical picture and model of SNR 1987A.

Accepted for publication in MNRAS Available from arXiv:1005.1037

A radial velocity and Calcium Triplet abundance survey of field Small Magellanic Cloud giants

R. De Propris^{1}, R.M. Rich^{2}, R. Mallery^{2} and C.D. Howard^{3}

1CTIO, Chile 2UCLA, USA 3SOFIA-USRA, USA

We present the results of a pilot wide-field radial velocity and metal abundance survey of red giants in 10 fields in the Small Magellanic Cloud (SMC). The targets lie at projected distances of 0.9 and 1.9 kpc from the SMC center (m − M = 18.79) to the north, east, south, and west. Two more fields are to the east at distances of 3.9 and 5.1 kpc. In this last field, we find only a few to no SMC giants, suggesting that the edge of the SMC in this direction lies approximately at 6 kpc from its center. In all eastern fields, we observe a double peak in the radial velocities of stars, with a component at the classical SMC recession velocity of ∼ 160 km s^{−1} and a high-velocity component at about 200 km s^{−1}, similar to observations in H i. In the most distant field (3.9 kpc), the low-velocity component is at 106 km s^{−1}. The metal abundance distribution in all fields is broad and centered at about [Fe/H] ∼ −1.25, reaching to solar and possibly slightly supersolar values and down to [Fe/H] of about −2.5. In the two innermost (0.9 kpc) northern and southern fields, we observe a secondary peak at metallicities of about ∼ −0.6. This may be evidence of a second episode of star formation in the center, possibly triggered by the interactions that created the Stream and Bridge.

Published in Astrophysical Journal Letters Available from arXiv:1003.6114

Photometry of two unusual A supergiant systems in the Small Magellanic Cloud

R.E. Mennickent^{1}, M.A. Smith^{2}, Z. Kołaczkowski^{1,3}, G. Pietrzyński^{4} and I. Soszyński^{4}

1Departamento de Astronomía, Facultad de Ciencias Físicas y Matemáticas, Universidad de Concepción, Casilla 160-C, Concepción, Chile 2Department of Physics, Catholic University of America, Washington, DC 20064, USA; Present address: Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218, USA 3Instytut Astronomiczny Uniwersytetu Wroclawskiego, Kopernika 11, 51-622 Wroclaw, Poland 4Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

We present multiwavelength broadband photometry and V, I time resolved photometry for two variable bright stars in the SMC, OGLE 004336.91−732637.7 (SMC-SC3) and OGLE 004633.76−731204.3 (SMC-SC4). The light curves span 12 years and show long-term periodicities (SMC-SC3) and modulated eclipses (SMC-SC4) that are discussed in terms of wide-orbit intermediate mass interacting binaries and associated envelopes. SMC-SC3 shows a primary period of 238.1 days along with a complicated waveform suggesting ellipsoidal variablity influenced by an eccentric orbit. This star also shows a secondary variability with an unstable periodicity that has a mean value of 15.3 days. We suggest this could be associated with nonradial pulsations.

Accepted for publication in MNRAS Available from arXiv:1004.2728

High dispersion spectroscopy of two A supergiant systems in the Small Magellanic Cloud with novel properties

R.E. Mennickent^{1} and M.A. Smith^{2}

1Universidad de Concepción, Departamento de Astronomía, Casilla 160-C, Concepción, Chile 2Department of Physics, Catholic University of America, Washington, DC 20064, USA

We present the results of a spectroscopic investigation of two novel variable bright blue stars in the SMC, OGLE 004336.91−732637.7 (SMC-SC3) and the periodically occulted star OGLE 004633.76−731204.3 (SMC-SC4), whose photometric properties were reported by Mennickent et al. (2010). High-resolution spectra in the optical and far-UV show that both objects are actually A + B type binaries. Three spectra of SMC-SC4 show radial velocity variations, consistent with the photometric period of 184.26 days found in Mennickent et al. 2010. The optical spectra of the metallic lines in both systems show combined absorption and emission components that imply that they are formed in a flattened envelope. A comparison of the radial velocity variations in SMC-SC4 and the separation of the V and R emission components in the Hα emission profile indicate that this envelope, and probably also the envelope around SMC-SC3, is a circumbinary disk with a characteristic orbital radius some three times the radius of the binary system. The optical spectra of SMC-SC3 and SMC-SC4 show, respectively, He i emission lines and discrete Blue Absorption Components (“BACs”) in metallic lines. The high excitations of the He i lines in the SMC-SC3 spectrum and the complicated variations of Fe ii emission and absorption components with orbital phase in the spectrum of SMC-SC4 suggests that shocks occur between the winds and various static regions of the stars’ co-rotating binary-disk complexes. We suggest that BACs arise from wind shocks from the A star impacting the circumbinary disk and a stream of former wind-efflux from the B star accreting onto the A star. The latter picture is broadly similar to mass transfer occurring in the more evolved (but less massive) Algol (B/A + K) systems, except that we envision transfer occurring in the other direction and not through the inner Lagrangian point. Accordingly, we dub these objects prototype of a small group of Magellanic Cloud wind-interacting A + B binaries.

Accepted for publication in MNRAS Available from arXiv:1005.1230

Chemical composition of the old globular clusters NGC 1786, NGC 2210 and NGC 2257 in the Large Magellanic Cloud.

A. Mucciarelli ^{1}, L. Origlia^{2} and F.R. Ferraro^{1}

1Dipartimento di Astronomia, Università di Bologna, Italy 2INAF — Osservatorio Astronomico di Bologna, Italy

This paper presents the chemical abundance analysis of a sample of 18 giant stars in 3 old globular clusters in the Large Magellanic Cloud, namely NGC 1786, NGC 2210 and NGC 2257. The derived iron content is [Fe/H]= −1.75 ± 0.01 dex (σ = 0.02 dex), −1.65 ± 0.02 dex (σ = 0.04 dex) and −1.95 ± 0.02 dex (σ = 0.04 dex) for NGC 1786, NGC 2210 and NGC 2257, respectively. All the clusters exhibit similar abundance ratios, with enhanced values (∼ +0.30 dex) of [α/Fe], consistent with the Galactic Halo stars, thus indicating that these clusters have formed from a gas enriched by Type II SNe. We also found evidence that r-process are the main channel of production of the measured neutron capture elements (Y, Ba, La, Nd, Ce and Eu). In particular the quite large enhancement of [Eu/Fe] (∼ +0.70 dex) found in these old clusters clearly indicates a relevant efficiency of the r-process mechanism in the LMC environment.

Accepted for publication in ApJ Available from arXiv:1005.0964

Testing mass loss in Large Magellanic Cloud Cepheids using infrared and optical observations II. Predictions and tests of the OGLE-III

fundamental-mode Cepheids

Hilding R. Neilson^{1,2}, Chow-Choong Ngeow^{3}, Shashi Kanbur^{4} and John B. Lester^{2,5}

1Argelander Institute of Astronomy, University of Bonn, Germany 2University of Toronto, Canada 3National Central University, Taiwan 4SUNY Oswego, USA 5University of Toronto Mississauga, Canada

In this article, we test the hypothesis that Cepheids have infrared excesses due to mass loss. We fit a model using the mass-loss rate and the stellar radius as free parameters to optical observations from the OGLE-III survey and infrared observations from the 2MASS and SAGE data sets. The sample of Cepheids have predicted minimum mass-loss rates ranging from zero to 10^{−8} M_{⊙} yr^{−1}, where the rates depend on the chosen dust properties. We use the predicted radii to compute the Period–Radius relation for LMC Cepheids, and to estimate the uncertainty caused by the presence of infrared excess for determining angular diameters with the infrared surface brightness technique. Finally, we calculate the linear and non-linear Period–Luminosity (P–L) relations for the LMC Cepheids at VIJHK + IRAC wavelengths and we find that the P–L relations are consistent with being non-linear at infrared wavelengths, contrary to previous results.

Accepted for publication in The Astrophysical Journal Available from arXiv:1005.0622

An XMM-Newton view of Planetary Nebulae in the Small Magellanic Cloud. The X-ray luminous central star of SMP SMC 22

S. Mereghetti^{1}, N, Krachmalnicoff^{1,5}, N. La Palombara^{1}, A. Tiengo^{1}, T. Rauch^{2}, F. Haberl^{3}, M.D. Filipović^{4} and R. Sturm^{3}

1INAF — IASF Milano, Italy 2IAA, Tübingen, Germany 3MPE, Garching, Germany 4University of Western Sydney, Australia 5University of Milano, Italy

During an X-ray survey of the Small Magellanic Cloud, carried out with the XMM-Newton satellite, we detected significant soft X-ray emission from the central star of the high-excitation planetary nebula SMP SMC 22. Its very soft spectrum is well fit with a non local thermodynamical equilibrium model atmosphere composed of H, He, C, N, and O, with abundances equal to those inferred from studies of its nebular lines. The derived effective temperature of 1.5 × 10^{5} K is in good agreement with that found from the optical/UV data. The unabsorbed flux in the 0.1–0.5 keV range is about 3 × 10^{−11} erg cm^{−2} s^{−1}, corresponding to a luminosity of 1.2 × 10^{37} erg/s at the distance of 60 kpc. We also searched for X-ray emission from a large number of SMC planetary nebulae,confirming the previous detection of SMP SMC 25 with a luminosity of (0.2–6)×10^{35} erg s^{−1} (0.1–1 keV). For the remaining objects that were not detected, we derived flux upper limits corresponding to luminosity values from several tens to hundreds times smaller than that of SMP SMC 22. The exceptionally high X-ray luminosity of SMP SMC 22 is probably due to the high mass of its central star, quickly evolving toward the white dwarf’s cooling branch, and to a small intrinsic absorption in the nebula itself.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1005.1031

An X-ray upper limit on the presence of a Neutron Star for the Small Magellanic Cloud and Supernova Remnant 1E 0102.2−7219

Michael J. Rutkowski^{1}, Eric M. Schlegel^{2}, Jonathan W. Keohane^{3} and Rogier A. Windhorst^{1}

1Arizona State University, USA 2University of Texas — San Antonio, USA 3Hampden Sydney College, USA

We present Chandra X-ray Observatory archival observations of the supernova remnant 1E 0102.2−7219, a young oxygen-rich remnant in the Small Magellanic Cloud. Combining 28 ObsIDs for 324 ks of total exposure time, we present an Advanced CCD Imaging Spectrometer image with an unprecedented signal-to-noise ratio (mean S/N ∼ \sqrt{S} = 6; maximum S/N > 35). We search within the remnant, using the source detection software wavdetect, for point sources which may indicate a compact object. Despite finding numerous detections of high significance in both broad and narrowband images of the remnant, we are unable to satisfactorily distinguish whether these detections correspond to emission from a compact object. We also present upper limits to the luminosity of an obscured compact stellar object which were derived from an analysis of spectra extracted from the high signal-to-noise image. We are able to further constrain the characteristics of a potential neutron star for this remnant with the results of the analysis presented here, though we cannot confirm the existence of such an object for this remnant.

Published in Astrophysical Journal Available from arXiv:1005.0635

The Optical Gravitational Lensing Experiment. The OGLE-III Catalog of Variable Stars. VIII. Type II Cepheids in the Small Magellanic Cloud

I. Soszyński^{1}, A. Udalski^{1}, M. K. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, Ł. Wyrzykowski^{3}, K. Ulaczyk^{1} and R. Poleski^{1}

1Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2Universidad de Concepción, Departamento de Astronomia, Casilla 160-C, Concepción, Chile 3Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK

The eighth part of the OGLE-III Catalog of Variable Stars (OIII-CVS) contains type II Cepheids in the Small Magellanic Cloud (SMC). The sample consists of 43 objects, including 17 BL Her, 17 W Vir and 9 RV Tau stars (first examples ever found in the SMC). Seven stars have been classified as peculiar W Vir stars — a recently identified subclass of type II Cepheids. These stars have distinctive light curves, are brighter and bluer than the ordinary W Vir variables. We confirm that a large fraction of the peculiar W Vir stars are members of binary systems. Three type II Cepheids exhibit eclipsing variations superimposed on the pulsation light curves, and three other objects show long-period ellipsoidal variability. All stars with the indication of binarity display secondary periods which may be interpreted as amplitude and/or phase modulations of the pulsation light curves with periods equal to the orbital periods or half the orbital periods. We do not have any model for these modulations, however this phenomenon rules out a possibility of the optical blends of a pulsating star and a binary system. For each object the multi-epoch V- and I-band photometry collected over 8 or 13 years of observations and finding charts are available to the astronomical community from the OGLE Internet archive.

Submitted to Acta Astronomica Available from arXiv:1005.3544

The youngest protostars in the Large Magellanic Cloud

M. Sewiło^{1}, R. Indebetouw^{2}, L. Carlson^{3}, B.A. Whitney^{4}, C.-H.R. Chen^{2}, M. Meixner^{1}, T. Robitaille^{5}, J.Th. van Loon^{6}, J.M. Oliveira^{6}, E. Churchwell^{7}, J.D. Simon^{8}, S. Hony^{9}, P. Panuzzo^{9}, M. Sauvage^{9}, J. Roman-Duval^{1}, K.

Gordon^{1}, C. Engelbracht^{10}, K. Misselt^{10}, K. Okumura^{9}, T. Beck^{1}, J. Hora^{5} and P.M. Woods^{11}

1Space Telescope Science Institute, USA 2University of Virginia, USA 3Johns Hopkins University, USA 4Space Science Institute, USA 5Harvard-Smithsonian Center for Astrophysics, USA 6School of Physical & Geographical Sciences, Lennard-Jones Laboratories, Keele University, UK 7University of Wisconsin — Madison, USA 8Observatories of the Carnegie Institution of Washington, USA 9CEA, Laboratoire AIM, France 10Steward Observatory, University of Arizona, USA 11Jodrell Bank Centre for Astrophysics, Alan Turing Building, School of Physics & Astronomy, University of Manchester, UK

We demonstrate the unique capabilities of Herschel to study very young luminous extragalactic young stellar objects (YSOs) by analyzing a central strip of the Large Magellanic Cloud obtained through the HERITAGE Science Demonstration Program. We combine PACS 100 and 160, and SPIRE 250, 350, and 500 μm photometry with 2MASS (1.25–2.17 μm) and Spitzer IRAC and MIPS (3.6–70 μm) to construct complete spectral energy distributions (SEDs) of compact sources. From these, we identify 207 candidate embedded YSOs in the observed region, ∼ 40% never before identified.We discuss their position in far-infrared color–magnitude space, comparing with previously studied, spectroscopically confirmed YSOs and maser emission. All have red colors indicating massive cool envelopes and great youth. We analyze four example YSOs, determining their physical properties by fitting their SEDs with radiative transfer models. Fitting full SEDs including the Herschel data requires us to increase the size and mass of envelopes included in the models. This implies higher accretion rates (greater or equal to 0.0001 M_{⊙} yr^{−1}), in agreement with previous outflow studies of high-mass protostars. Our results show that Herschel provides reliable longwave SEDs of large samples of high-mass YSOs; discovers the youngest YSOs whose SEDs peak in Herschel bands; and constrains the physical properties and evolutionary stages of YSOs more precisely than was previously possible.

Accepted for publication in A&A Letters (Herschel Special Issue)

A detailed look at chemical abundances in Magellanic Cloud Planetary Nebulae. I. The Small Magellanic Cloud

Richard A. Shaw^{1}, Ting-Hui Lee^{1,2}, Letizia Stanghellini^{1}, James E. Davies^{1}, D.A. García-Hernández^{3,4}, Pedro García-Lario^{5}, José V. Perea-Calderón^{6}, Eva Villaver^{7}, Arturo Manchado^{3,4,8}, Stacy Palen^{9} and Bruce Balick^{10}

1National Optical Astronomy Observatory, Tucson, AZ, USA 2Department of Physics & Astronomy, Western Kentucky University, Bowling Green, KY, USA 3Instituto de Astrofísica de Canarias, vía Láctea s/n, La Laguna, E-38200 Tenerife, Spain; also CSIC, Spain 4Departamento de Astrofísica, Universidad de La Laguna (ULL), E-38205 La Laguna, Tenerife, Spain 5Herschel Science Centre, European Space Astronomy Centre, Research and Scientific Support Department of ESA, Villafranca del Castillo, P.O. Box 50727 E-28080 Madrid, Spain 6European Space Astronomy Centre, INSA S.A., P.O. Box 50727 E-28080 Madrid, Spain 7Universidad Autónoma de Madrid, Departamento de Física Teórica C-XI, 28049 Madrid, Spain 89 Consejo Superior de Investigaciones Cientificas, E-28006 Madrid, Spain 9Physics Department, Weber State University, Ogden, UT, USA 10Astronomy Department, Univ. of Washington, Seattle, WA, USA

We present an analysis of elemental abundances of He, N, O, Ne, S, and Ar in Magellanic Cloud planetary nebulae (PNe), and focus initially on 14 PNe in the Small Magellanic Cloud (SMC). We derived the abundances from a combination of deep, high dispersion optical spectra, as well as mid-infrared (IR) spectra from the Spitzer Space Telescope. A detailed comparison with prior SMC PN studies shows that significant variations among authors of relative emission line flux determinations lead to systematic discrepancies in derived elemental abundances between studies that are > ∼0.15 dex, in spite of similar analysis methods. We used ionic abundances derived from IR emission lines, including those from ionization stages not observable in the optical, to examine the accuracy of some commonly used recipes for ionization correction factors (ICFs). These ICFs, which were developed for ions observed in the optical and ultraviolet, relate ionic abundances to total elemental abundances. We find that most of these ICFs work very well even in the limit of substantially sub-Solar metallicities, except for PNe with very high ionization. Our abundance analysis shows enhancements of He and N that are predicted from prior dredge-up processes of the progenitors on the AGB, as well as the well known correlations among O, Ne, S, and Ar that are little affected by nucleosynthesis in this mass range. We identified MG 8 as an interesting limiting case of a PN central star with a ∼ 3.5 M_{⊙} progenitor in which hot-bottom burning did not occur in its prior AGB evolution. We find no evidence for O depletion in the progenitor AGB stars via the O-N cycle, which is consistent with predictions for lower-mass stars. We also find low S/O ratios relative to SMC H ii regions, with a deficit comparable to what has been found for Galactic PNe. Finally, the elemental abundances of one object, SMP-SMC 11, are more typical of SMC H ii regions, which raises some doubt about its classification as a PN.

Accepted for publication in Astrophysical Journal Available from arXiv:1005.2221

Cold dust clumps in dynamically hot gas

S. Kim^{1}, E. Kwon^{1}, S.C. Madden^{2}, M. Meixner^{3,19}, S. Hony^{2}, P. Panuzzo^{2}, M. Sauvage^{2}, J. Roman-Duval^{3}, K.D. Gordon^{3}, C. Engelbracht^{4}, F.P. Israel^{5}, K. Misselt^{4}, K. Okumura^{2}, A. Li^{6}, A. Bolatto^{7}, R. Skibba^{4}, F. Galliano^{2},

M. Matsuura^{8,9}, J.-P. Bernard^{10}, C. Bot^{11}, M. Galametz^{2}, A. Hughes^{12,13}, A. Kawamura^{14}, T. Onishi^{15}, D.

Paradis^{16}, A. Poglitsch^{17}, W.T. Reach^{16,18}, T. Robitaille^{19}, M. Rubio^{20} and A.G.G.M. Tielens^{5}

1Astronomy & Space Science, Sejong University, 143-747, Seoul, South Korea 2CEA Laboratoire AIM, Irfu/SAp, Orme des Merisiers, F-91191 Gif-sur-Yvette, France 3Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 4Steward Observatory, University of Arizona, 933 North Cherry Ave., Tucson, AZ 85721, USA 5Sterrewacht Leiden, Leiden University, P.O. Box 9513, NL-2300 RA Leiden, The Netherlands 6314 Physics Building, Department of Physics and Astronomy, University of Missouri-Columbia, Columbia, MO 65211, USA 7Department of Astronomy, Lab for Millimeter-wave Astronomy, University of Maryland, College Park, MD 20742-2421, USA 8Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 9Mullard Space Science Laboratory, University College London, Holmbury St. Mary, Dorking, Surrey RH5 6NT, United Kingdom 10Centre d’ Étude Spatiale des Rayonnements, CNRS, 9 av. du Colonel Roche, BP 4346, 31028 Toulouse, France 11Observatoire Astronomique de Strasbourg, 11, rue de l’université, 67000 Strasbourg, France 12Centre for Supercomputing and Astrophysics, Swinburne University of Technology, Hawthorn VIC 3122, Australia 13CSIRO Australia Telescope National Facility, 76 Epping Rd., NSW1710, Australia 14Department of Astrophysics, Nagoya University, Chikusa-ku, Nagoya 464-8602 , Japan 15Department of Physical Science, Osaka Prefecture University, Gakuen 1-1, Sakai, Osaka 599-8531, Japan 16Spitzer Science Center, California Institute of Technology, MS 220-6, Pasadena, CA 91125, USA 17Max-Planck-Institut für extraterrestrische Physik, Gießenbachstraße, 85748 Garching, Germany 18Stratospheric Observatory for Infrared Astronomy, Universities Space Research Association, Mail Stop 211-3, Moffett Field, CA 94035, USA 19Center for Astrophysics, 60 Garden St., MS 67 , Harvard University, Cambridge, MA 02138, USA 20Departamento de Astronomia, Universidad de Chile, Casilla 36-D, Santiago, Chile

Aims. We present clumps of dust emission from Herschel observations of the Large Magellanic Cloud (LMC) and their physical and statistical properties. We catalog cloud features seen in the dust emission from Herschel observations of the LMC, the Magellanic type irregular galaxy closest to the Milky Way, and compare these features with H i catalogs from the ATCA+Parkes H i survey. Methods. Using an automated cloud-finding algorithm, we identify clouds and clumps of dust emission and examine the cumulative mass distribution of the detected dust clouds. The mass of cold dust is determined from physical parameters that we derive by performing spectral energy distribution fits to 250, 350, and 500 μm emission from SPIRE observations using dust grain size distributions for graphite/silicate in low-metallicity extragalactic environments. Results. The dust cloud mass spectrum follows a power law distribution with an exponent of γ = −1.8 for clumps larger than 4 × 10^{2} M_{⊙} and is similar to the H i mass distribution. This is expected from the theory of ISM structure in the vicinity of star formation.

Accepted for publication in A&A Letters (Herschel Special Issue) Available from arXiv:1005.3123

VLT multi-object spectroscopy of 33 eclipsing binaries in the Small

Magellanic Cloud. New distance and depth of the SMC, and a record-breaking apsidal motion

P. North^{1}, R. Gauderon^{1}, F. Barblan^{2} and F. Royer^{3}

1Laboratoire d’Astrophysique, École Polytechnique Fédérale de Lausanne (EPFL), Switzerland 2Geneva Observatory, Geneva University, Switzerland 3GEPI, UMR 8111 du CNRS, Observatoire de Paris-Meudon, France

Our purpose is to provide reliable stellar parameters for a significant sample of eclipsing binaries, which are representative of a whole dwarf and metal-poor galaxy. We also aim at providing a new estimate of the mean distance to the SMC and of its depth along the line of sight for the observed field of view. We use radial velocity curves obtained with the ESO FLAMES facility at the VLT and light curves from the OGLE-II photometric survey. The radial velocities were obtained by least-squares fits of the observed spectra to synthetic ones, excluding the hydrogen Balmer lines. Our sample contains 23 detached, 9 semi-detached and 1 overcontact systems. Most detached systems have properties consistent with stellar evolution calculations from single-star models at the standard SMC metallicity Z = 0.004, though they tend to be slightly overluminous. The few exceptions are probably due to third light contribution or insufficient signal-to-noise ratio. The mass ratios are consistent with a flat distribution, both for detached and semi-detached/contact binaries. A mass-luminosity relation valid from ∼ 4 to ∼ 18 M_{⊙} is derived. The uncertainties are in the ±2 to ±11% range for the masses, in the ±2 to ±5% range for the radii and in the ±1 to ±6% range for the effective temperatures. The average distance modulus is 19.11 ± 0.03 (66.4 ± 0.9 kpc). The moduli derived from the V and from the I data are consistent within 0.01 mag. The 2-σ depth of the SMC is, for our field, of 0.25 mag or 7.6 kpc under the assumption of a gaussian distribution of stars along the line of sight. Three systems show significant apsidal motion, one of them with an apsidal period of 7.6 years, the shortest known to date for a detached system with main sequence stars.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1005.4336

Neutral material around the B[e] supergiant star LHA 115-S 65. An outflowing disk or a detached Keplerian rotating disk?

M. Kraus^{1}, M. Borges Fernandes^{2,3} and F.X. de Araújo^{3}

1Astronomický ústav, Akademie věd České republiky, Fričova 298, 251 65 Ondřejov, Czech Republic 2UMR 6525 H. Fizeau, Univ. Nice Sophia Antipolis, CNRS, Observatoire de la Côte d’Azur, Av. Copernic, F-06130 Grasse, France 3Observatório Nacional, Rua General José Cristino 77, 20921-400 São Cristovão, Rio de Janeiro, Brazil

B[e] supergiants are surrounded by large amounts of hydrogen neutral material, traced by the emission in the optical [O i] lines. This neutral material is most plausibly located within their dense, cool circumstellar disks, which are formed from the (probably non-spherically symmetric) wind material released by the star. Neither the formation mechanism nor the resulting structure and internal kinematics of these disks (or disk-like outflows) are well known. However, rapid rotation, lifting the material from the equatorial surface region, seems to play a fundamental role. The B[e] supergiant LHA 115-S 65 (in short: S 65) in the Small Magellanic Cloud is one of the two most rapidly rotating B[e] stars known. Its almost edge-on orientation allows a detailed kinematical study of its optically thin forbidden emission lines. With a focus on the rather strong [O i] lines, we intend to test the two plausible disk scenarios: the outflowing and the Keplerian rotating disk. Based on high- and low-resolution optical spectra, we investigate the density and temperature structure in those disk regions that are traced by the [O i] emission to constrain the disk sizes and mass fluxes needed to explain the observed [O i] line luminosities. In addition, we compute the emerging line profiles expected for either an outflowing disk or a Keplerian rotating disk, which can directly be compared to the observed profiles. Both disk scenarios deliver reasonably good fits to the line luminosities and profiles of the [O i] lines. Nevertheless, the Keplerian disk model seems to be the more realistic one, because it also agrees with the kinematics derived from the large number of additional lines in the spectrum. As additional support for the presence of a high-density, gaseous disk, the spectrum shows two very intense and clearly double-peaked [Ca ii] lines. We discuss a possible disk-formation mechanism, and similarities between S 65 and the group of Luminous Blue Variables.

Accepted for publication in A&A Available from arXiv:1005.4511

Dust in the bright supernova remnant N 49 in the LMC

M. Otsuka^{1}, J.Th. van Loon^{2}, K.S. Long^{1}, M. Meixner^{1}, M. Matsuura^{3}, W.T. Reach^{4,5}, J. Roman-Duval^{1}, K. Gordon^{1}, M. Sauvage^{6}, S. Hony^{6}, K. Misselt^{7}, C. Engelbracht^{7}, P. Panuzzo^{6}, K. Okumura^{6}, P.M. Woods^{8}, F.

Kemper^{8} and G. Sloan^{9}

1Space Telescope Science Institute, USA 2School of Physical & Geographical Sciences, Lennard-Jones Laboratories, Keele University, UK 3Department of Physics and Astronomy, University College London, UK 4Spitzer Science Center, California Institute of Technology, USA 5Stratospheric Observatory for Infrared Astronomy, Universities Space Research Association, USA 6CEA, Laboratoire AIM, Irfu/SAp, France 7Steward Observatory, University of Arizona, USA 8Jodrell Bank Centre for Astrophysics, UK 9Department of Astronomy, Cornell University, USA

We investigate the dust associated with the supernova remnant (SNR) N 49 in the Large Magellanic Cloud (LMC) as observed with the Herschel Space Observatory. N 49 is unusually bright because of an interaction with a molecular cloud along its eastern edge. We have used PACS and SPIRE to measure the far IR flux densities of the entire SNR and of a bright region on the eastern edge of the SNR where the SNR shock is encountering the molecular cloud. Using these fluxes supplemented with archival data at shorter wavelengths, we estimate the dust mass associated with N 49 to be about 10 M_{⊙}. The bulk of the dust in our simple two-component model has a temperature of 20–30 K, similar to that of nearby molecular clouds. Unfortunately, as a result of the limited angular resolution of Herschel at the wavelengths sampled with SPIRE, the uncertainties are fairly large. Assuming this estimate of the dust mass associated with the SNR is approximately correct, it is probable that most of the dust in the SNR arises from regions where the shock speed is too low to produce significant X-ray emission. The total amount of warm 50–60 K dust is ∼ 0.1 or 0.4 M_{⊙}, depending on whether the dust is modeled in terms of carbonaceous or silicate grains. This provides a firm lower limit to the amount of shock heated dust in N 49.

Accepted for publication in Astronomy & Astrophysics Letters (Herschel Special Issue) Available from arXiv:1005.2787

Thesis

Supernova remnants as a probe of dust grains in the interstellar medium

Brian J. Williams^{1}

1North Carolina State University, USA

Interstellar dust grains play a crucial role in the evolution of the galactic interstellar medium (ISM). Despite its importance, however, dust remains poorly understood in terms of its origin, composition, and abundance throughout the universe. Supernova remnants (SNRs) provide a laboratory for studying the evolution of dust grains, as they are one of the only environments in the universe where it is possible to observe grains being both created and destroyed. SNRs exhibit collisionally heated dust, allowing dust to serve as a diagnostic both for grain physics and for the plasma conditions in the SNR. I present theoretical models of collisionally heated dust which calculate grain emission as well as destruction rates. In these models, I incorporate physics such as nonthermal sputtering caused by grain motions through the gas, a more realistic approach to sputtering for small grains, and arbitrary grain compositions porous and composite grains. I apply these models to infrared and X-ray observations of Kepler’s supernova and the Cygnus Loop in the galaxy, and SNRs 0509−67.5, 0519−69.0, and 0540−69.3 in the LMC. X-ray observations characterize the hot plasma while IR observations constrain grain properties and destruction rates. Such a multi-wavelength approach is crucial for a complete understanding of gas and dust interaction and evolution. Modeling of both X-ray and IR spectra allows disentangling of parameters such as pre and postshock gas density, as well as swept-up masses of gas and dust, and can provide constraints on the shock compression ratio. Observations also show that the dust-to-gas mass ratio in the ISM is lower by a factor of several than what is inferred by extinction studies of starlight. Future observatories, such as the James Webb Space Telescope and the International X-ray Observatory, will allow testing of models far beyond what is possible now.

Ph.D. Thesis Available from arXiv:1005.1296 and from http://www.lib.ncsu.edu/theses/available/etd-03092010-171032/